you would like to save php 10,000 for to your debut birthday party.how much should you deposit in q sqvings account every end of the month for 3 years if the interest is at 0.12% compounded monthy?
1
Expert's answer
2021-10-08T15:49:30-0400

Amount needed to be saved, A is 10000

Interest rate, r is 0.12% compounded monthly

Time, t is 3 years

Let the amount that needs to be deposited be P

So;

A"=" P"[" 1"+\frac{r}{n}]^{nt}"

P"[" 1"+\frac{0.12}{1200}]^{12}\times^{3}" "=" 10000

P"[" 1"+0.0001]^{36}" "=" 10000

1.0001"^{36}" P"=" 10000

1.003606P"=" 10000

P is 9964.069

Answer is Php9964.06